If the distance between the earth and the sun were half its present value, the number of days in a year would have been
(a) 64.5
(b) 129
(c) 182.5
(d) 730

Grade:6

1 Answers

Deepak Patra
9 years ago

(b) According to Kepler’s law T21 / T22 = R31­ / R32

Here T1 = 365 days ; T2 = ?; R1 = R; R2 = R/2

? T2 = T1 (R2 / R1)3/2 = 365 [R/2 / R]3/2 = 129 days